# Revisiting DDIM Inversion for Controlling Defect Generation by Disentangling the Background

### v1

@@ Abstract
In anomaly detection, the scarcity of anomalous data compared to normal data poses a challenge in effectively utilizing deep neural network representations to identify anomalous features. From a data-centric perspective, generative models can solve this data imbalance issue by synthesizing anomaly datasets. Although previous research tried to enhance the controllability and quality of generating defects, they do not consider the relation between background and defect. Since the defect depends on the object’s background (i.e., the normal part of an object), training only the defect area cannot utilize the background information, and even generation can be biased depending on the mask information. In addition, controlling logical anomalies should consider the dependency between background and defect areas (e.g., orange colored defect on a orange juice bottle). In this paper, our paper proposes modeling a relationship between the background and defect, where background affects denoising defects; however, the reverse is not. We introduce the regularizing term to disentangle denoising background from defects. From the disentanglement loss, we rethink defect generation with DDIM Inversion, where we generate the defect on the target normal image. Additionally, we theoretically prove that our methodology can generate a defect on the target normal image with an invariant background. We demonstrate our synthetic data is realistic and effective in several experiments.
[Uncaptioned image][Uncaptioned image]
Figure 1: Left image is a comparison between ours and baselines for MVTec AD. Right image is a comparison between ours and baselines for MVTec Loco
@@ 1 Introduction
Data scarcity presents a significant challenge in industrial anomaly detection, as imbalanced datasets can lead to overfitting in deep neural networks, damaging inspection performance. Limited labeled training examples introduce model bias, preventing the detection model from effectively generalizing to unseen anomalies. Recent unsupervised anomaly detection approaches [23, 5, 17] mitigate the imbalance issue by avoiding reliance on anomaly instances during training. However, these methods may struggle when the backbone model experiences distribution shifts in feature space, making it difficult to distinguish normal and anomalous features. Additionally, they often fall short in detecting fine-grained anomalies at the pixel level.
@@
Recent research on defect generation aims to address data scarcity through a data-centric approach, synthesizing anomaly data to train on augmented samples. GAN-based methods [6] generate realistic synthetic anomaly data from limited examples, improving model performance on rare anomalies; however, they lack control over defect location, often transforming normal regions into anomalies unintentionally. Diffusion-based methods [10] improve control by specifying defect locations and blending normal latent backgrounds with synthetic defect latents. While these personalization techniques enable defect placement on object backgrounds, some challenges remain unaddressed.
First, focusing solely on controllability with masks can compromise the quality of generation. Since the defect’s appearance is dependent on the background of the target image, using a misleading mask may result in the generation of a biased anomaly that does not align with the data distribution. For example, defects should not appear outside the objects themselves. Additionally, defects should vary based on the texture of the background and the location of the target mask such as Figure 3. Generating logical anomalies, in particular, becomes challenging without accounting for the relationship between the defect and the background, especially when compared to generating structural anomalies, which are easier to control.
Secondly, overfitting to few-shot anomaly instances can lead to biased generation, where controllable defect generation may fail when encountering unseen backgrounds. Since anomaly instances consist of both foreground (defect) and background, it is crucial to model the relationship between these two components to mitigate overfitting issues. This consideration helps ensure that the model can generate defects effectively, avoiding the transfer of inadequate feature of the defect.
@@
To address the issues mentioned above, we propose modeling the relationship between the defect and background to achieve both high-quality and controllable defect generation. Our work suggests the disentanglement loss function for generating faithful defects on the target normal instances. Our formulation ensures independence from defect to background, allowing the denoising process for defects to proceed without affecting the background’s denoising process. At the same time, the background influences the defect area’s denoising process, enhancing defect generation by reducing reliance on the mask information. In addition, we propose the inference strategy to generate defect on the target normal latent, where we apply DDIM Inversion to the normal latent and generate defects on the background. We theoretically show that we can initialize anomaly latent by noising the target area of inversed normal latent while reconstructing its background. To sum up, we make the following contributions.
@@ 2.1 Industrial anomaly detection
Industrial anomaly detection (IAD) has seen significant progress, transitioning from traditional methods reliant on labeled data to more advanced machine learning and computer vision approaches. By leveraging pretrained models to extract normal features, unsupervised anomaly detection methods focus on identifying anomalous regions in products [4, 3, 23, 5, 17]. Memory-based approaches [4, 3, 23] store normal feature representations and calculate similarity scores to pinpoint anomalous regions. However, since unsupervised methods do not incorporate anomalous features during training, accurately detecting defect areas in anomaly instances remains challenging. As a result, supervised or semi-supervised learning methods remain highly effective for identifying fine-grained defect areas at the pixel level, especially when sufficient anomaly instances are available for training.
@@ 2.2 Defect generation
Collecting sufficient anomalous data is challenging due to the time-consuming nature and high annotation costs involved. To address this scarcity, data augmentation and generation methods have been developed. Cut-Paste [14] employs pixel-level data augmentation, while Crop-Paste [16] focuses on feature-level augmentation to help the model learn discriminative features between normal and anomalous data. However, since these augmentation-based methods do not model distributions of anomalies, generating a faithful defect dataset remains difficult.
In the DFMGAN [6] paper, the authors leverage Style-GAN [12] to generate defects from a limited number of anomalous data. While they also incorporate normal data to guide defect generation, their approach lacks the ability to control defect generation on unseen target normal data. In contrast, AnomalyDiffusion [10] addresses this by training a Text-to-Image diffusion model to control defect generation on target normal images. However, their method does not train the diffusion process to account for the background of anomalous data, meaning that defect generation does not consider the relationship between the defect and the target normal background. Since some defects are dependent on the background, it is crucial to model the interaction between the background and defect to generate more faithful and contextually accurate defects.
@@ 2.3 Diffusion model
Since the advent of diffusion models [9, 26], text-to-image (T2I) models [20] that control image generation through text has been developed. The denoising network $\epsilon_{\theta}$ is conditioned on text embedding $C$, where text embedding controls the output with cross-attention. Latent variable, $z_{t}$ is injected by Gaussian noise $\epsilon$ from the $z_{0}$ as follows [9]:
$z_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,\ where\ \epsilon\sim\mathcal{N}(0,I)$ (1)
@@
$min_{\theta}E_{z_{0},\epsilon\sim\mathcal{N}(0,I),t\sim Uniform(1,T)}||\epsilon-\epsilon_{\theta}(z_{t},t,C)||_{2}^{2}$ (2)
@@ 2.4 DDIM Inversion
By the DDIM Sampling [25], we can denoise the noise with non-Markovian deterministic diffusion process as follows:
@@
Especially, from the DDIM Inversion [25], editing of target images is enabled with changing text embeddings [7, 18]. Furthermore, target images can be translated to source image concepts via controlling cross-attention. In this paper, we utilize the DDIM Inversion to personalize the target normal images and generate the defect in the target area.
@@ 3 Method
3.1 Loss formulation
As discussed earlier, our objective is to model the relationship between defect and background in order to generate more accurate defect representations on normal images. In the U-Net architecture, attention mechanisms facilitate interaction between defect and background regions, effectively reducing random noise. While defects may vary depending on the background context, the background itself should remain unaffected by defect generation. To decouple the influence of defects on the background, we propose the following loss formulation with masked noising strategy.
@@ 3.1 Loss formulation
In Eq.3.1, minimizing $Matching\ Loss$ denoises that defect area of $z_{t}$ with considering the noised background area. Since $(1-m)\odot z_{t}=z_{t}^{m}$ is satisfied for all $t$, minimizing $Regularizer$ term denoises the background area of $z_{t}$, where the denoising process of the background area of $z_{t}$ is independent on the defect area of $z_{t}$. We prove this statement in Lemma 1.
Lemma 1.
$(1-m)\odot z_{0}=(1-m)\odot\sqrt{1\over\alpha_{t}}[z_{t}^{m}-\sqrt{1-\alpha_{t}}\epsilon_{\theta^{*}}(z_{t}^{m},t,C_{2})]$
According to Lemma 1, it is possible to reconstruct the background without requiring information about the defect area. As a result, the model can produce a variety of defects while generating the same background. The proof of Lemma 1 is provided in Appendix.
@@ 3.2 DDIM Inversion for generating synthetic data
After reversing the latent, $\tilde{z}_{T}^{m}$, the denoising path should generate the defect on the mask area. To guarantee defect generation on the target mask, we present the following theorem to demonstrate that the standard initialization strategy works. We prove Theorem 1 in Appendix.
Theorem 1.
@@
Suppose that Theorem 1, we randomly initialized the latent, $z_{T}^{*}=\tilde{z}_{T}^{m}+m\odot\epsilon$, where $\epsilon\sim\mathcal{N}(0,I)$. From Eq.3.2, denoising $z_{t}^{*}$ generates the defect on the mask area affected by the denoised background part of latent, $z_{t}^{m}$.
@@
Although we initialize the masked area with random noise, Proposition 1 proves that denoising $z_{t}^{*}$ maintains the DDIM-Inversion gap, where we kept the background of the target normal image.
Proposition 1.
@@
From Proposition 1, defect generation on target mask still does not affect the denoising of background after DDIM Inversion. We prove Proposition 1 in Appendix.
@@ 3.3.1 Masking attention strategy
As previously discussed, our objective is to control defect generation under the disentangled background, where its denosing process is independent of another area of latent. In the Text-to-Image (T2I) model, U-Net is designed with self-attention and cross-attention mechanisms, where cross-attention facilitates interaction between text embeddings and latent features.
To disentangle each text embedding, we define masked cross-attentions of $z_{t}$ and $z_{t}^{m}$ inspired by [19]. In the process of cross-attention, $m^{\prime}$ is resized mask and each attention map is defined as follows: $M:=[M_{def},M_{bg}]=[softmax({QK_{def}^{T}\over\tau}),softmax({QK_{bg}^{T}\over\tau})]$ and $M^{m}:=[M_{m}^{m},M_{bg}^{m}]=[softmax({Q^{m}K_{m}^{T}\over\tau}),softmax({Q^{m}K_{bg}^{T}\over\tau})]$, where $Q$ and $Q^{m}$ are from query vectors from $z_{t}$ and $z_{t}^{m}$, respectively. $K_{def}$, $K_{bg}$, and $K_{m}$ are the key vectors from the text encoder embeddings $C_{def}$, $C_{bg}$, and $C_{m}$, respectively. Then, the masked cross-attention is given by
@@
Here, $\hat{Q}^{m}$, $\hat{K}^{m}$, $\hat{V}^{m}$ are the query, key, values on the self-attention layer from the masked latent $z_{t}^{m}$.
Since $(1-m)\odot\epsilon_{\theta}(z_{t}^{m},t,C_{2})$ denoises that the background of $z_{t}$, Eq.3.3.1 shows that its denoising step only depends on the background of $z_{t}$. The details will be given in Supplementary
@@ 3.3.2 Refining target mask with attention
During inference, we guide defect generation within the target mask $m$; however, ts generation may not be filled within the target mask. This occurs because background latents can influence the defect region, occasionally incorporating background elements into the generated defect. To address this, we refine the target mask to ensure a more precise defect localization. This refined mask is then used to construct a robust training dataset for anomaly detection.
@@ 4.1 Experimental settings
Dataset  In this paper, we conduct our experiments on industrial anomaly detection benchmark datasets: MVTec-AD [1] and MVTec-Loco [2]. Each dataset includes anomaly instances in the test set, while the training set contains only normal instances. To train the generative model on these anomaly instances, we split each anomaly dataset into two folds: one fold is used for training the generative model, and the other for evaluating inspection performance in anomaly detection. For unsupervised settings, we exclude anomaly instances from training, as unsupervised methods do not learn from anomaly data.
Details of experiments  We utilize the Stable Diffusion [21] for fine-tuning anomaly dataset with 500 iterations, where each model is trained on the defect type of objects. We only trained the U-Net part of Stable Diffusion followed by Dreambooth [24] and inferred the sample with DDIM sampler [25] with 50-time steps.
Baseline  We compared our methodology with other defect generation-based methodologies; DFMGAN [6] and Anomalydiffusion [10], where each setting is identical in their papers. In addition, we compared ours with Unsupervised-based anomaly detection; RD4AD [5], Patchcore [23], and SimpleNet [17].

### v2

# Background-Aware Defect Generation for Robust Industrial Anomaly Detection

@@ Abstract
Detecting anomalies in industrial settings is challenging due to the scarcity of labeled anomalous data. Generative models can mitigate this issue by synthesizing realistic defect samples, but existing approaches often fail to model the crucial interplay between defects and their background. This oversight leads to unrealistic anomalies, especially in scenarios where contextual consistency is essential (i.e., logical anomaly). To address this, we propose a novel background-aware defect generation framework, where the background influences defect denoising without affecting the background itself by ensuring realistic synthesis while preserving structural integrity. Our method leverages a disentanglement loss to separate the background’s denoising process from the defect, enabling controlled defect synthesis through DDIM Inversion. We theoretically demonstrate that our approach maintains background fidelity while generating contextually accurate defects. Extensive experiments on MVTec-AD and MVTec-Loco benchmarks validate our method’s superiority over existing techniques in both defect generation quality and anomaly detection performance.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Left image is a comparison between ours and baselines for MVTec AD. Right image is a comparison between ours and baselines for MVTec Loco.
@@
Data scarcity presents a significant challenge in industrial anomaly detection, as imbalanced datasets can lead to overfitting in deep neural networks, damaging inspection performance. Recent unsupervised anomaly detection approaches (Roth et al., 2022; Deng & Li, 2022; Liu et al., 2023) mitigate the imbalance issue by avoiding reliance on anomaly instances during training. However, these methods may struggle when the backbone model experiences distribution shifts in feature space, making it difficult to distinguish normal and anomalous features. Additionally, they often fall short in detecting fine-grained anomalies at the pixel level.
@@
Recent research on defect generation aims to address data scarcity through a data-centric approach, synthesizing anomaly data to train on augmented samples. GAN-based methods (Duan et al., 2023) generate realistic synthetic anomaly data from limited examples, improving model performance on rare anomalies; however, they lack control over defect location, often transforming normal regions into anomalies unintentionally. Diffusion-based methods (Hu et al., 2024) improve control by specifying defect locations and blending normal latent background with synthetic defect latent. While these methods have made strides in generating faithful anomaly samples, they fail to adequately model the critical relationship between the defect and its background.
Defects are closely related to the background, making it difficult to control their generation using only mask information. If a model relies solely on a given mask, the generated defect may appear out of place, failing to integrate naturally with the object’s structure. Even when the defect is generated within the masked region, it can overfit to the mask shape rather than capturing realistic variations, leading to unrealistic anomalies in Figure 1, where defect is not aligned with the background. Such inconsistencies in generated data can ultimately degrade the robustness of anomaly detection models.
Furthermore, in the case of logical anomalies, defects are not always visually apparent, making it necessary to leverage background information to detect them effectively. Since these anomalies are defined not just by their appearance but by their contextual inconsistency, modeling the relationship between the defect and its background is essential for generating meaningful and realistic anomalies, ultimately enhancing the robustness of the detection model.
While modeling the relationship between defects and their background is crucial, it is essential to structure this relationship so that the background influences the defect rather than the reverse. Achieving controllability in defect generation requires the model to generate defects directly on target normal samples, where defect is not contained. However, training a generative model with anomaly samples presents a challenge: defects and background are inherently entangled in these examples, making it difficult for the model to separate their contributions. By decoupling the background from the defect during training, model can generate realistic defects that align with the characteristics of the background in the target normal sample.
@@
Therefore, our work suggests the disentanglement loss function for generating faithful defects on the target normal instances. Our formulation ensures independence from defect to background, allowing the denoising process for defects to proceed without affecting the background’s denoising process. At the same time, the background influences the defect area’s denoising process, enhancing defect generation by reducing reliance on the mask information. In addition, we propose the inference strategy to generate defect on the target normal latent, where we apply DDIM Inversion to the normal latent and generate defects on the background. We theoretically show that we can initialize anomaly latent by noising the target area of inversed normal latent while reconstructing its background. To sum up, we make the following contributions.
@@ 2.1 Industrial anomaly detection
Industrial anomaly detection (IAD) has seen significant progress, transitioning from traditional methods reliant on labeled data to more advanced machine learning and computer vision approaches. By leveraging pretrained models to extract normal features, unsupervised anomaly detection methods focus on identifying anomalous regions in products (Defard et al., 2021; Cohen & Hoshen, 2020; Roth et al., 2022; Deng & Li, 2022; Liu et al., 2023). Memory-based approaches (Defard et al., 2021; Cohen & Hoshen, 2020; Roth et al., 2022) store normal feature representations and calculate similarity scores to pinpoint anomalous regions. However, since unsupervised methods do not incorporate anomalous features during training, accurately detecting defect areas in anomaly instances remains challenging. As a result, supervised or semi-supervised learning methods remain highly effective for identifying fine-grained defect areas at the pixel level, especially when sufficient anomaly instances are available for training.
@@ 2.2 Defect generation
Collecting sufficient anomalous data is challenging due to the time-consuming nature and high annotation costs involved. To address this scarcity, data augmentation and generation methods have been developed. Cut-Paste (Li et al., 2021) employs pixel-level data augmentation, while Crop-Paste (Lin et al., 2020) focuses on feature-level augmentation to help the model learn discriminative features between normal and anomalous data. However, since these augmentation-based methods do not model distributions of anomalies, generating a faithful defect dataset remains difficult.
In the DFMGAN (Duan et al., 2023) paper, the authors leverage Style-GAN (Karras et al., 2020) to generate defects from a limited number of anomalous data. While they also incorporate normal data to guide defect generation, their approach lacks the ability to control defect generation on unseen target normal data. In contrast, AnomalyDiffusion (Hu et al., 2024) addresses this by training a Text-to-Image diffusion model to control defect generation on target normal images. However, their method does not train the diffusion process to account for the background of anomalous data, meaning that defect generation does not consider the relationship between the defect and the target normal background. Since some defects are dependent on the background, it is crucial to model the interaction between the background and defect to generate more faithful and contextually accurate defects.
Refer to captionRefer to caption
@@ 2.3 Diffusion model
Since the advent of diffusion models (Ho et al., 2020; Song et al., 2021b), text-to-image (T2I) models (Rombach et al., 2022a) that control image generation through text has been developed. The denoising network $\epsilon_{\theta}$ is conditioned on text embedding $C$, where text embedding controls the output with cross-attention. Latent variable, $z_{t}$ is injected by Gaussian noise $\epsilon$ from the $z_{0}$ as follows (Ho et al., 2020):
$z_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,\epsilon\sim\mathcal{N}(0,I)$ (1)
@@
$\min_{\theta}E_{z_{0},\epsilon\sim\mathcal{N}(0,I),t\sim Uniform(1,T)}||\epsilon-\epsilon_{\theta}(z_{t},t,C)||_{2}^{2}$ (2)
@@ 2.4 DDIM Inversion
By the DDIM Sampling (Song et al., 2021a), we can denoise the noise with non-Markovian deterministic diffusion process as follows:
@@
Especially, from the DDIM Inversion (Song et al., 2021a), editing of target images is enabled with changing text embeddings (Hertz et al., 2023; Mokady et al., 2023). Furthermore, target images can be translated to source image concepts via controlling cross-attention. In this paper, we utilize the DDIM Inversion to personalize the target normal images and generate the defect in the target area.
@@ 3 Method
3.1 Disentanglement Loss formulation
As discussed earlier, our objective is to model the relationship between defect and background in order to generate more accurate defect representations on normal images. In the U-Net architecture, attention mechanisms facilitate interaction between defect and background regions, effectively reducing random noise. While defects may vary depending on the background context, the background itself should remain unaffected by defect generation. To decouple the influence of defects on the background, we propose a disentanglement loss that ensures defect generation is conditioned on the background while preventing interference with background reconstruction.
@@ 3.1 Disentanglement Loss formulation
In Eq.3.1, minimizing $Matching\ Loss$ denoises that defect area of $z_{t}$ with considering the noised background area. Since $(1-m)\odot z_{t}=z_{t}^{m}$ is satisfied for all $t$, minimizing $Regularizer$ term denoises the background area of $z_{t}$, where the denoising process of the background area of $z_{t}$ is independent on the defect area of $z_{t}$. We prove this statement in Lemma 3.1.
Lemma 3.1.
Suppose that $\mathcal{L}(\theta^{*})=0$, then background of $z_{t}$ is reconstructed as follows: $(1-m)\odot z_{0}=(1-m)\odot\sqrt{1\over\alpha_{t}}[z_{t}^{m}-\sqrt{1-\alpha_{t}}\epsilon_{\theta^{*}}(z_{t}^{m},t,C_{2})]$
According to Lemma 3.1, it is possible to reconstruct the background without requiring information about the defect area. As a result, the model can produce a variety of defects while generating the same background. The proof of Lemma 3.1 is provided in Appendix A.2.
@@ 3.2 DDIM Inversion for generating synthetic data
After reversing the latent, $\tilde{z}_{T}^{m}$, the denoising path should generate the defect on the mask area. To guarantee defect generation on the target mask, we present the following theorem to demonstrate that the standard initialization strategy works. We prove Theorem 3.2 in Appendix A.3.
Theorem 3.2.
@@
Suppose that Theorem 3.2, we randomly initialized the latent, $z_{T}^{*}=\tilde{z}_{T}^{m}+m\odot\epsilon$, where $\epsilon\sim\mathcal{N}(0,I)$. From Eq.3.2, denoising $z_{t}^{*}$ generates the defect on the mask area affected by the denoised background part of latent, $z_{t}^{m}$.
@@
Although we initialize the masked area with random noise, Proposition 3.3 proves that denoising $z_{t}^{*}$ maintains the DDIM-Inversion gap, where we kept the background of the target normal image.
Proposition 3.3.
@@
From Proposition 3.3, defect generation on target mask still does not affect the denoising of background after DDIM Inversion. We prove Proposition 3.3 in Appendix A.4.
@@ 3.3.1 Masking attention strategy
In the Text-to-Image (T2I) model, U-Net is designed with self-attention and cross-attention mechanisms, where cross-attention facilitates interaction between text embeddings and latent features.
To disentangle each text embedding, we define masked cross-attentions of $z_{t}$ and $z_{t}^{m}$ inspired by (Park et al., 2024). In the process of cross-attention, $m^{\prime}$ is resized mask and each attention map is defined as follows: $M:=[M_{def},M_{bg}]=[softmax({QK_{def}^{T}\over\tau}),softmax({QK_{bg}^{T}\over\tau})]$ and $M^{m}:=[M_{m}^{m},M_{bg}^{m}]=[softmax({Q^{m}K_{m}^{T}\over\tau}),softmax({Q^{m}K_{bg}^{T}\over\tau})]$, where $Q$ and $Q^{m}$ are from query vectors from $z_{t}$ and $z_{t}^{m}$, respectively. $K_{def}$, $K_{bg}$, and $K_{m}$ are the key vectors from the text encoder embeddings $C_{def}$, $C_{bg}$, and $C_{m}$, respectively. Then, the masked cross-attention is given by
@@
Here, $\hat{Q}^{m}$, $\hat{K}^{m}$, $\hat{V}^{m}$ are the query, key, values on the self-attention layer from the masked latent $z_{t}^{m}$. Since $(1-m)\odot\epsilon_{\theta}(z_{t}^{m},t,C_{2})$ denoises that the background of $z_{t}$, Eq.3.3.1 shows that its denoising step only depends on the background of $z_{t}$. The details will be given in Appendix A.5.
@@ 3.3.2 Refining target mask with attention
During inference, we guide defect generation within the target mask $m$; however, its generation may not be filled within the target mask. This occurs because background latents can influence the defect region, occasionally incorporating background elements into the generated defect. To address this, we refine the target mask to ensure a more precise defect localization. This refined mask is then used to construct a robust training dataset for anomaly detection.
@@ 4.1 Experimental settings
Dataset  In this paper, we conduct our experiments on industrial anomaly detection benchmark datasets: MVTec-AD (Bergmann et al., 2019) and MVTec-Loco (Bergmann et al., 2022). Each dataset includes anomaly instances in the test set, while the training set contains only normal instances. To train the generative model on these anomaly instances, we split each anomaly dataset into two folds: one fold is used for training the generative model, and the other for evaluating inspection performance in anomaly detection. For unsupervised settings, we exclude anomaly instances from training, as unsupervised methods do not learn from anomaly data.
Details of experiments  We utilize the Stable Diffusion (Rombach et al., 2022b) for fine-tuning anomaly dataset with 500 iterations, where each model is trained on the defect type of objects. We only trained the U-Net part of Stable Diffusion followed by Dreambooth (Ruiz et al., 2023) and inferred the sample with DDIM sampler (Song et al., 2021a) with 50-time steps.
Baseline  We compared our methodology with other defect generation-based methodologies; DFMGAN (Duan et al., 2023) and Anomalydiffusion (Hu et al., 2024), where each setting is identical in their papers. In addition, we compared ours with Unsupervised-based anomaly detection; RD4AD (Deng & Li, 2022), Patchcore (Roth et al., 2022), and SimpleNet (Liu et al., 2023).